OpenStudy (animecore1):

I will metal for help and explanation Marty made a $220 bank deposit using $10 bills and $5 bills. She gave the teller a total of 38 bills, how many $5 bills were in the deposit? A) 6 five-dollar bills B) 28 five-dollar bills C) 32 five-dollar bills D) 34 five-dollar bills

OpenStudy (anonymous):

hmm its imposible two know exacly i mean mabey a or b i dont quite know also i saw an admin with 100 smartscore

OpenStudy (animecore1):

I figured it out, 32*5=160 that means there are 6 bills left which are 10's 10*6=60 160+60=220 Therefor it is C can't believe I couldn't figure it out lol

OpenStudy (anonymous):

what grade is this

OpenStudy (anonymous):

darn i ask that alot

OpenStudy (animecore1):

8th